# Task Arithmetic Through The Lens Of One-Shot Federated Learning

### v1

@@ 5.1.2 Merging without training heterogeneity
In Section 3.2, we analyzed how training heterogeneity causes objective inconsistency, degrading Task Arithmetic’s performance. To validate this, we compare its performance on task vectors fine-tuned via homogeneous and heterogeneous training. While in the experiments conducted by [28], all task vectors were fine-tuned using a consistent learning rate $1\mathrm{e}{-5}$ and 2000 iterations for homogeneous fine-tuning, our approach in Section 5.1.1 employs heterogeneous fine-tuning. We compare the performance of Task Arithmetic on these two sets of task vectors to validate our theoretical findings.
Table 3 summarizes the performance of Task Arithmetic with heterogeneous fine-tuning and with homogeneous fine-tuning in the experiment of merging all eight task vectors. Again we report the average normalized accuracy and the normalized accuracy for each dataset. As evident from the table, Task Arithmetic with homogeneous fine-tuning consistently outperforms its heterogeneous counterpart across all datasets, except for SUN397.
Table 4 and Figure 2 compare Task Arithmetic’s performance under homogeneous and heterogeneous fine-tuning across 247 experiments. Homogeneous fine-tuning outperforms heterogeneous fine-tuning in $92.31\%$ of cases, as shown in Table 4. Moreover, Figure 2 shows that homogeneous fine-tuning can improve the average normalized accuracy by up to more than $30\%$. These results highlight the significant negative impact of training heterogeneity on the performance of Task Arithmetic.
@@ Methods
Heterogenous Fine-Tuning
@@
Homogenous Fine-Tuning
@@
Table 3: Using Task Arithmetic to combine eight task vectors from heterogeneous and homogeneous fine-tuning processes. Each method is evaluated on eight datasets, with normalized accuracy reported for each.
@@
Table 4: Percentage of improved, unchanged, and degraded experiments using task vectors with homogeneous fine-tuning process compared to those with heterogeneous fine-tuning process. The method used to combine task vectors is Task Arithmetic.
Figure 2: Histogram showing the change in average normalized accuracy when using task vectors from homogeneous fine-tuning compared to heterogeneous fine-tuning. A positive value on x-axis indicates that the performance of Task Arithmetic using homogeneous fine-tuned task vectors is better than those of using heterogeneous fine-tuned task vectors, while a negative value indicates the opposite. The y-axis represents the number of experiments within the range of change values.
@@ 5.2 Discussion on Experimental Results
While Section 3.1 highlights how data heterogeneity degrades Task Arithmetic, our experimental results in Section 5.1 show it is less problematic compared to training heterogeneity. First of all, FedNova, designed to address training heterogeneity, consistently outperforms Task Arithmetic more frequently and significantly than FedGMA, which targets data heterogeneity. As shown in Table 2 and Figure 2, FedNova not only enhances the merged model’s performance more frequently but also yields greater overall performance gains than FedGMA.
Second, among Federated Learning algorithms, CCLIP and Median demonstrate the best performance. As discussed in Section 4, these methods are designed for robust aggregation in the presence of outliers. In our setting, they effectively address training heterogeneity which causes certain task vectors to have disproportionately large norms and behave like outliers. For example, the cross-validation process selects a much larger learning rate of $1\mathrm{e}{-4}$ for SVHN, compared to $1\mathrm{e}{-5}$ used for other datasets. This hyperparameter setup results in the SVHN task vector having a significantly larger norm (reported in Appendix C.1.5), making it an outlier that negatively impacts the merged model’s performance on other tasks when using Task Arithmetic. By employing robust aggregation methods like Median and CCLIP, we reduce the influence of the SVHN task vector, which improves the merged model’s performance on other tasks.
@@
Further evidence is presented in Appendix C.2, where we evaluate the performance of Median, FedGMA and CCLIP on task vectors generated through homogeneous fine-tuning. Using the performance of Task Arithmetic on these homogeneously fine-tuned task vectors as the baseline, we find that Federated Learning algorithms rarely improve upon the baseline. In fact, Task Arithmetic consistently emerges as the best-performing approach when merging these task vectors generated without training heterogeneity. This reinforces our observation that training heterogeneity is a more significant issue than data heterogeneity in practice.
@@ 6 Experiments on Merging LLMs
We now present and discuss our experimental results on merging LLMs for three tasks: instruction following, mathematical reasoning, and code generation. We follow the experimental paradigm of [86]. We merge task vectors constructed by three models—WizardLM-13B [79], WizardMath-13B [48], and Llama-2-13B-Code-Alpaca [5]—for instruction following, mathematical reasoning, and code generation, respectively. All three models are fine-tuned from Llama2-13B [67]. For instruction following, we evaluate the models on AlpacaEval [45]. For mathematical reasoning, we use GSM8K [10] and MATH [27]. For code generation, we evaluate on HumanEval [7] and MBPP [2]. Performance metrics include win rate for AlpacaEval, zero-shot accuracy for GSM8K and MATH, and pass@1 for HumanEval and MBPP.
Since all models used in this experiment are downloaded from HuggingFace, we do not have access to their fine-tuning hyperparameter settings. As a result, FedNova cannot be applied in this experiment because it requires knowledge of learning rates and the number of iterations, which are unavailable. Furthermore, when implementing Median, taking the median of two vectors reduces to averaging, which is equivalent to Task Arithmetic. Consequently, we implement Median only for merging three task vectors, with the corresponding results deferred to Appendix D.2. For additional details on experiments, please refer to Appendix D.
@@ 6.1 Experimental Results
In Table 5, we compare the performance of three methods: Task Arithmetic, FedGMA, and CCLIP. The results show that when merging two out of three task vectors, FedGMA and CCLIP often outperform Task Arithmetic. However, when merging all three task vectors, Task Arithmetic demonstrates superior performance on code generation and instruction-following tasks. Notably, Task Arithmetic consistently excels in instruction-following tasks, achieving either the highest accuracy or accuracy comparable to the other methods.
@@ 6.2 Discussion on Experimental Results
In Section 3.2, we theoretically analyzed how using different learning rates and number of iterations creates training heterogeneity and thus leads to objective inconsistency. However, our experimental results in Section 6 reveal that employing different fine-tuning methods further exacerbates training heterogeneity. For instance, in our experiments, the Llama-2-13B-Code-Alpaca model is fine-tuned using QLoRA [11], a parameter-efficient fine-tuning (PEFT) approach. PEFT adjusts only a small subset of parameters while leaving the rest unchanged [25]. Consequently, task vectors generated by PEFT typically have smaller norms compared to those generated through standard fine-tuning. This discrepancy can pose challenges when merging task vectors. Simply regulating the behavior of task vectors with larger norms can lead to unintended negative effects, and, to date, no Federated Learning algorithm has been specifically designed to address this issue.
@@
In our experiments, Llama-2-13B-Code-Alpaca, which is fine-tuned for code generation using PEFT, produces a task vector with a notably small norm of 5.05. In contrast, WizardLM-13B and WizardMath-13B, fine-tuned for instruction following and mathematical reasoning via standard fine-tuning, generate task vectors with much larger norms of 142.61 and 52.62, respectively. This significant disparity in task vector norms between code generation and instruction following leads to complications when merging task vectors by using Federated Learning algorithms. As shown in Table 5, when merging tasks include both instruction following (which has a large norm) and code generation (which has a small norm), FedGMA and CCLIP either fail to outperform Task Arithmetic or achieve comparable performance on these two tasks. This highlights that addressing training heterogeneity by focusing solely on differences in hyperparameters is insufficient in practice.
@@
While some studies have explored the challenges of merging large models fine-tuned via PEFT [87, 77], merging PEFT-generated task vectors with those produced by standard fine-tuning remains an open research question. Further investigation is required to devise effective strategies for combining such task vectors in Task Arithmetic. Additionally, more research is needed to develop robust aggregation methods in Federated Learning to address this type of practical training heterogeneity.
7 Related Work
As our work bridges Federated Learning and Task Arithmetic, a prominent approach within the growing domain of model merging, this section reviews related work on both model merging and Federated Learning.
7.1 Model Merging
Task Arithmetic is one of many recent works on model merging [74, 18, 75, 51, 1, 12]. Though the term model merging is relatively new, firstly formalized by [51], the concept has received significant investigation [75, 74, 29, 82, 65, 24]. For example, [75] averages the pre-trained model parameters and fine-tuned parameters to enhance the robustness of fine-tuned model against distribution shifts. [74] averages parameters of multiple fine-tuned models with different hyperparameter configurations can improve robustness and accuracy. [29] averages several parameters along the same trajectory of SGD can lead to better generalization.
Task Arithmetic, introduced by [28], refines model merging by introducing task vectors and a hyperparameter $\lambda$ to control how much task vectors modify pre-trained model parameters. This method has inspired various follow-up work on using simple arithmetic operations for model merging [86, 87, 55, 63, 80, 63, 81] such as sparsifying task vectors [86], merging parameter-efficient modules [87], fine-tuning in linearized model spaces [55] and resolving sign interference of task vectors [80]. A concurrent work of Task Arithmetic is [31], who propose the RegMean, inspired by the process of merging linear regression models. They also note that model merging is an extreme case of Federated Learning, where only a single round of communication occurs. This aligns with the core idea of our work, where we view model merging as a form of one-shot Federated Learning. However, our work delves deeper into this notion, providing more detailed explanations and analysis.
There is a substantial body of research dedicated to understanding the effectiveness of model merging [13, 15, 19, 20, 3, 39, 17, 60]. Some studies focus on the theory of linear model connectivity [13, 15, 19, 20], while others emphasize the flatness of the loss landscape [3, 39, 17, 60]. However, there has been relatively little work addressing the effectiveness of Task Arithmetic except for [55, 57].
7.2 Federated Learning
In Federated Learning, FedAvg [52] is widely used to solve the following distributed optimization problem across $M$ devices
$\min_{\theta\in\mathbb{R}^{d}}L(\theta):=\frac{1}{M}\sum_{m=1}^{M}L_{m}(\theta)$ (16)
where $L_{m}(\theta):=\mathbb{E}_{x_{m}\sim\mathcal{D}_{m}}[\ell(\theta,x_{m})]$ is the objective function on each device $m$, defined by some loss function $\ell$ and data distribution $\mathcal{D}_{m}$. The core idea behind FedAvg is to perform local SGD on each device, followed by model averaging on the server. There is a substantial body of research analyzing the performance of FedAvg and local SGD [43, 34, 35, 36, 73, 71, 21, 56, 68, 72].
A key challenge in the theoretical analysis of FedAvg arises from data heterogeneity, where each device $m$ has a different data distribution $\mathcal{D}_{m}$. In the homogeneous setting where $\mathcal{D}_{m}=\mathcal{D}\;\forall m$, [72] has established the min-max complexity of FedAvg with smooth and convex loss functions. In the more complex heterogeneous setting, various works have derived the convergence rate of FedAvg under different assumptions about data heterogeneity [43, 35, 36, 73, 71, 21, 56, 68]. In this work, rather than focusing on extending existing theoretical results, we focus on using these results to analyze Task Arithmetic.
To address the challenges posed by data heterogeneity, extensive research has focused on designing algorithms to improve the performance of FedAvg for Federated Learning [34, 59, 42, 84, 44, 66, 69]. Some work has enhanced optimization algorithms by regulating the local training process [34, 59, 42, 69, 44], while other papers have proposed alternative aggregation methods beyond simple averaging [66, 84]. Another line of research focuses on personalized Federated Learning [61, 64, 50, 41, 16], addressing data heterogeneity by adapting the global model locally for each device.
Aside from data heterogeneity, [69] notice different local training process (which we refer to training heterogeneity) exacerbates the convergence of federated optimization algorithms, leading them to converge to a stationary point of an objective function inconsistent with the original objective function.
8 Conclusions
@@
In this paper, we establish a connection between Task Arithmetic and one-shot Federated Learning. By leveraging theoretical insights from Federated Learning, we identify and analyze two key sources of heterogeneity—data heterogeneity and training heterogeneity—and their impact on Task Arithmetic. Also, we adapt Federated Learning algorithms, demonstrating their great potentials to significantly improve the performance of Task Arithmetic for model merging. We hope this work serves as a foundation for advancing the understanding, enhancing algorithms, and expanding the applications of Task Arithmetic through the lens of Federated Learning.
@@ Acknowledgments
We would like to thank Kasper Vinken and Mehdi Bahrami for insightful discussions and valueful feedback for the project.
Contribution Statement
@@ Appendix A Task Arithmetic Property
In this section, we review a paper that provides theoretical insights into Task Arithmetic [55]. We examine the relationship between data heterogeneity and the Task Arithmetic property proposed in their work. To facilitate a stronger connection between their framework and our perspective, we adapt their definition of the Task Arithmetic property as follows.
@@
(Task Arithmetic Property 1 from [55]) Consider a set of task vectors $\{\tau_{t}\}_{t=1}^{T}$ with associated task data distributions $\{\mathcal{D}_{t}\}_{t=1}^{T}$. Suppose all the distributions $\{\mathcal{D}_{t}\}_{t=1}^{T}$ have non-intersecting supports. Let $f$ be a network function. We say a network function $f$ satisfies the Task Arithmetic property around $\theta_{0}$ with respect to $\{\tau_{t}\}_{t=1}^{T}$ and $\{\mathcal{D}_{t}\}_{t=1}^{T}$ if
@@
(Necessary Condition) Suppose the Task Arithmetic property holds true and all the objective functions $L_{t}$ are $H$-smooth and convex in $\theta$. Further assume $\theta_{0}+\tau_{t}$ is optimal for $L_{t}$, i.e., $\nabla_{\theta}L_{t}(\theta_{0}+\tau_{t})=0$. Then $\theta_{0}+\lambda\sum_{t=1}^{T}\tau_{t}$ is optimal for $L$. Moreover, the data heterogeneity at $\theta_{0}+\lambda\sum_{t=1}^{T}\tau_{t}$ is zero, i.e.,
@@
The above proposition shows that data heterogeneity being zero is actually a necessary condition for optimal performance of Task Arithmetic. In other words, the parameters generated by Task Arithmetic have to be a shared optimum for all $L_{t}$.
Appendix B Challenges in Adapting Federated Learning Algorithms for Task Arithmetic
Selecting the right Federated Learning algorithms to implement requires a clear understanding of key challenges that complicate the adaptation. In this section, we analyze two key challenges.
First, the number of communication rounds is limited. As Task Arithmetic is only one-shot Federated Learning, algorithms relying on multiple communication rounds are unsuitable. For instance, some Federated Learning algorithms add regularization terms to local objective functions [42, 14] to encourage local updates to remain close to the global model parameters transmitted from the previous communication round. However, in our one-shot setting, only the pre-trained model parameters $\theta_{0}$ are communicated, so applying this type of regularization would constrain each task’s fine-tuned parameters to be near the pre-trained parameters, potentially degrading both convergence and task-specific performance.
Other algorithms, like those using variance reduction techniques [34, 59] or adaptively updating the server’s optimal outer step size [30, 46], aim to address heterogeneity through an iterative process. These iterative methods require either each local device or the central server to compute, accumulate, and update certain metrics over multiple communication rounds. Since Task Arithmetic operates within a one-shot setting, implementing such iterative updates is impossible. This constraint limits the use of these approaches to modify Task Arithmetic, as they cannot perform the necessary progressive adjustments over time.
Second, no additional training is allowed. In conventional Federated Learning, alternative aggregation methods can also be implemented at the server to counteract the effects of data heterogeneity. However, many of these approaches impose additional computational cost. For instance, [84, 76] ask each device to compute metrics comparing local and global data, which are then used as additional scores for aggregation.
Third, no additional datasets are available. Many Federated Learning algorithms rely on supplementary datasets, which is not feasible for modifying Task Arithmetic. In one-shot Federated Learning, a common approach to address data heterogeneity is knowledge distillation [88, 23, 40]. These methods often require access to extra datasets from which either local devices or the central server distills knowledge to improve model performance.
Given the constraints and unique needs for adapting Federated Learning algorithms, we propose the criteria for selecting Federated Learning algorithms as in Section 4.
Appendix C Merging CLIP
C.1 Additional Experiment Details on Merging CLIP

### v2

@@ 6.1.2 Merging without training heterogeneity
In Section 4.2, we analyzed how training heterogeneity causes objective inconsistency, degrading Task Arithmetic’s performance. While in the experiments conducted by Ilharco et al. (2023), all task vectors were homogeneously fine-tuned using a consistent learning rate $1\mathrm{e}{-5}$, our approach in Section 6.1.1 employs heterogeneous fine-tuning. We compare the performance of Task Arithmetic on these two sets of task vectors to validate our theoretical findings in Section 4.2.
Table 3 compares Task Arithmetic’s performance on heterogeneously and homogeneously fine-tuned task vectors when merging all eight task vectors. Again we report the average normalized accuracy and the normalized accuracy for each dataset. As evident from the table, Task Arithmetic with homogeneous fine-tuning consistently outperforms its heterogeneous counterpart across all datasets, except for SUN397.
Table 4 and Figure 2 compare Task Arithmetic’s performance on homogeneously and heterogeneously fine-tuned task vectors across 247 experiments. Homogeneous fine-tuning outperforms heterogeneous fine-tuning in $92.31\%$ of cases, as shown in Table 4. Moreover, Figure 2 shows that homogeneous fine-tuning can improve the average normalized accuracy by up to more than $30\%$. These results highlight the significant negative impact of training heterogeneity on the performance of Task Arithmetic.
@@ Methods
Heterogeneous Fine-Tuning
@@
Homogeneous Fine-Tuning
@@
Table 3: Using Task Arithmetic to merge eight task vectors from heterogeneous and homogeneous fine-tuning processes. Each method is evaluated on eight datasets, with normalized accuracy reported for each dataset.
@@
Table 4: Percentage of improved, unchanged, and degraded experiments using task vectors with homogeneous fine-tuning process compared to those with heterogeneous fine-tuning process.
Figure 2: Histogram showing the change in average normalized accuracy when using task vectors from homogeneous fine-tuning compared to heterogeneous fine-tuning. The x-axis shows the difference in average normalized accuracy relative to Task Arithmetic, with positive values indicating improvement and negative values indicating degradation. The y-axis represents the number of experiments within each range of change values.
@@ 6.2 Discussion on Experimental Results
While Section 4.1 highlights how data heterogeneity degrades Task Arithmetic, our experimental results in Section 6.1 show it is less problematic compared to training heterogeneity. First, FedNova, designed to address training heterogeneity, consistently outperforms Task Arithmetic more frequently and significantly than FedGMA which targets data heterogeneity. Table 2 and Figure 1 demonstrate that FedNova both improves performance more frequently and achieves greater overall gains than FedGMA.
Second, among Federated Learning algorithms, CCLIP and Median demonstrate the best performance. As discussed in Section 5, these methods are designed for robust aggregation in the presence of outliers. In our setting, they effectively address training heterogeneity which causes certain task vectors to have disproportionately large norms and behave like outliers. For example, the cross-validation process selects a much larger learning rate of $1\mathrm{e}{-4}$ for SVHN, compared to $1\mathrm{e}{-5}$ used for other datasets. This hyperparameter setup results in the SVHN task vector having a significantly larger norm (reported in Appendix B.1.5), making it an outlier that negatively impacts the merged model’s performance on other tasks when using Task Arithmetic. By employing robust aggregation methods like Median and CCLIP, we reduce the influence of the SVHN task vector, which improves the merged model’s performance on other tasks.
@@
Further evidence is presented in Appendix B.2, where we evaluate the performance of Median, FedGMA and CCLIP on task vectors generated via homogeneous fine-tuning. Using the performance of Task Arithmetic on these homogeneously fine-tuned task vectors as the baseline, we find that Federated Learning algorithms rarely improve upon the baseline. In fact, Task Arithmetic consistently emerges as the best-performing approach when merging those task vectors generated without training heterogeneity. This reinforces our observation that training heterogeneity is a more significant issue than data heterogeneity in practice.
@@ 7 Experiments on Merging LLMs
We now present and discuss our experimental results on merging LLMs for three tasks: instruction following, mathematical reasoning, and code generation. We follow the experimental paradigm of Yu et al. (2024). We merge task vectors from three models—WizardLM-13B (Xu et al., 2023), WizardMath-13B (Luo et al., 2023), and Llama-2-13B-Code-Alpaca (Chaudhary, 2023)—for instruction following, mathematical reasoning, and code generation, respectively. All three models are fine-tuned from Llama2-13B (Touvron et al., 2023). For instruction following, we evaluate the models on AlpacaEval (Li et al., 2023a). For mathematical reasoning, we use GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021). For code generation, we use HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021). Performance metrics include win rate for AlpacaEval, zero-shot accuracy for GSM8K and MATH, and pass@1 for HumanEval and MBPP.
Since all models used in this experiment are downloaded from HuggingFace, we do not have access to their fine-tuning hyperparameter settings. As a result, FedNova cannot be applied in this experiment because it requires knowledge of learning rates and the number of iterations, which are unavailable. Furthermore, when implementing Median, taking the median of two vectors reduces to averaging, which is equivalent to Task Arithmetic. Thus, we implement Median only for merging three task vectors, with the corresponding results deferred to Appendix C.2. For additional details on experiments, please refer to Appendix C.
@@ 7.1 Experimental Results
In Table 5, we compare the performance of three methods: Task Arithmetic, FedGMA, and CCLIP. The results show that when merging two out of three task vectors, FedGMA and CCLIP often outperform Task Arithmetic. However, when merging all three task vectors, Task Arithmetic shows superior performance on code generation and instruction-following tasks. Notably, Task Arithmetic consistently excels in instruction-following tasks, achieving either the highest accuracy or accuracy comparable to the other methods.
@@ 7.2 Discussion on Experimental Results
In Section 4.2, we theoretically analyzed how using different learning rates and number of iterations creates training heterogeneity and thus leads to objective inconsistency. However, our experimental results in Section 7 reveal that employing different fine-tuning methods further exacerbates training heterogeneity. Specifically, in our experiments, the Llama-2-13B-Code-Alpaca model is fine-tuned using QLoRA (Dettmers et al., 2024), a parameter-efficient fine-tuning (PEFT) approach.
@@
PEFT adjusts only a small subset of parameters while leaving the rest unchanged (Han et al., 2024). Consequently, task vectors generated by PEFT typically have smaller norms compared to those generated through standard fine-tuning. In our experiments, Llama-2-13B-Code-Alpaca, which is fine-tuned for code generation using PEFT, produces a task vector with a notably small norm of 5.05. In contrast, WizardLM-13B and WizardMath-13B, fine-tuned for instruction following and mathematical reasoning via standard fine-tuning, generate task vectors with much larger norms of 142.61 and 52.62, respectively. This discrepancy can pose challenges when merging task vectors. Simply regulating the behavior of task vectors with larger norms can lead to unintended negative effects. As shown in Table 5, when merging tasks include both instruction following (large norm) and code generation (small norm), FedGMA and CCLIP either fail to outperform Task Arithmetic or achieve comparable performance on these two tasks. This highlights that addressing training heterogeneity by focusing solely on differences in hyperparameters is insufficient in practice.
@@
While some studies have explored the challenges of merging large models fine-tuned via PEFT (Zhang et al., 2023; Wu et al., 2024), merging PEFT-generated task vectors with those produced by standard fine-tuning remains an open research question. Further investigation is required to design effective strategies for combining such task vectors in Task Arithmetic. Additionally, more research is needed to develop robust aggregation methods in Federated Learning to address this type of practical training heterogeneity.
8 Limitations and Conclusion
Limitations First, in Federated Learning, there remains a gap in theoretical understanding and algorithmic development for addressing data heterogeneity in large-scale experiments. This might arise from the lack of a universal notion of data heterogeneity. While existing theoretical work primarily focuses on first-order heterogeneity such as Assumption 4.5, recent studies (Patel et al., 2024; Wang et al., 2022) suggest that this perspective may be insufficient to fully capture data heterogeneity in practice. This underscores the need for more refined definitions of data heterogeneity and the development of corresponding algorithms.
Second, a recent study (Yadav et al., 2024b) shows that when merging very large models such as PaLM-2-64B (Anil et al., 2023), the differences in performance between various merging methods tend to diminish. While Section 4.1 discusses how a high-quality pre-trained model mitigates the adverse effects of data heterogeneity, our framework does not yet provide a theoretical explanation for why the large size of the pre-trained model can also alleviate data heterogeneity, and we leave this as an open question for future work.
@@ 8 Limitations and Conclusion
Conclusion In this paper, we established a connection between Task Arithmetic and one-shot Federated Learning. By leveraging theoretical insights from Federated Learning, we identified and analyzed two key sources of heterogeneity—data heterogeneity and training heterogeneity—and their impact on Task Arithmetic. Also, we adapted Federated Learning algorithms for model merging, demonstrating their great potential to significantly improve over the performance of Task Arithmetic. We hope this work serves as a foundation for advancing the understanding, enhancing the algorithms, and expanding the applications of Task Arithmetic through the lens of Federated Learning.
Author Contributions
@@ Acknowledgments
We thank the reviewers and action editor for their insightful comments and constructive feedback, which have significantly improved the quality of this work. We are also grateful to Kasper Vinken and Mehdi Bahrami for valuable discussions and feedback throughout the project.
@@ Appendix A Task Arithmetic Property
In this section, we review a paper that provides theoretical insights into Task Arithmetic (Ortiz-Jimenez et al., 2023). We examine the relationship between data heterogeneity and the Task Arithmetic property proposed in their work. To facilitate a stronger connection between their framework and our perspective, we adapt their definition of the Task Arithmetic property as follows.
@@
(Task Arithmetic Property 1 from Ortiz-Jimenez et al. (2023)) Consider a set of task vectors $\{\tau_{t}\}_{t=1}^{T}$ with associated task data distributions $\{\mathcal{D}_{t}\}_{t=1}^{T}$. Suppose all the distributions $\{\mathcal{D}_{t}\}_{t=1}^{T}$ have non-intersecting supports. Let $f$ be a network function. We say a network function $f$ satisfies the Task Arithmetic property around $\theta_{0}$ with respect to $\{\tau_{t}\}_{t=1}^{T}$ and $\{\mathcal{D}_{t}\}_{t=1}^{T}$ if
@@
(Necessary Condition) Suppose the Task Arithmetic property holds true and all the objective functions $L_{t}$ are $H$-smooth and convex in $\theta$. Further assume $\theta_{0}+\tau_{t}$ is optimal for $L_{t}$, i.e., $\nabla_{\theta}L_{t}(\theta_{0}+\tau_{t})=0$. Then $\theta_{0}+\lambda\sum_{t=1}^{T}\tau_{t}$ is optimal for $L$, and the data heterogeneity at $\theta_{0}+\lambda\sum_{t=1}^{T}\tau_{t}$ is zero, i.e.,
@@
The above proposition shows that data heterogeneity being zero is actually a necessary condition for the optimal performance of Task Arithmetic. In other words, the parameters generated by Task Arithmetic have to be a shared optimum for all $L_{t}$.
Appendix B Merging CLIP
B.1 Additional Experiment Details on Merging CLIP